STARK EFFECT USING TENSORIAL FORMALISM IN THE D_{2h} GROUP: APPLICATION TO THE C_{2}H_{4} MOLECULE

W. RABALLAND, \underline{M. ROTGER}, V. BOUDON, M. LOETE, Laboratoire de Physique de l’Université de Bourgogne, CNRS UMR 5027, 9, Avenue Alain Savary, BP 47870, F-21078 Dijon Cedex, France.

A tensorial formalism adapted to the case of X_{2}Y_{4} asymmetric molecules^{a} with D_{2h} symmetry has been developed in the same way as in the previous works on XY_{4}^{b} (T_{d}) and XY_{6}^{c} (O_{h}) spherical tops, XY_{5}Z (C_{4v}) symmetric tops^{d}^{,}^{e} or XY_{2}Z_{2} (C_{2v}) asymmetric tops^{f}. We have then constructed a Stark Hamiltonian using the same principle. This model allows the calculation of Stark shifts and splittings in the spectra of D_{2h} molecules. Preliminary predictions will be shown for some rovibrational bands of the C_{2}H_{4} molecule.

aW. Raballand, M. Rotger, V. Boudon and M. Loëte, J. Mol. Spectrosc., 217, 239-248, (2003). bJ.-P. Champion, M. Loëte and G. Pierre, in “Spectroscopy of the Earth’s Atmosphere and Interstellar Medium” (K. N. Rao and

A. Weber, Eds.) pp. 339-422, Academic Press, Inc., San Diego (1992).

cN. Cheblal, M. Loëte and V. Boudon, J. Mol. Spectrosc., 197, 222-231, (1999). dM. Rotger, V. Boudon and M. Loëte, J. Mol. Spectrosc., 200, 123-130, (2000). eM. Rotger, V. Boudon and M. Loëte, J. Mol. Spectrosc., 200, 131-137, (2000). fM. Rotger, V. Boudon and M. Loëte, J. Mol. Spectrosc., 216, 297-307, (2002).